# Quantitative Understanding of PDF Fits and their Uncertainties

### v1

@@ 1 Introduction
The complex nature of inverse problems has prompted the development of sophisticated statistical methods and tools to tackle them. In general, PDF determinations can be broadly classified into two main categories, depending on whether a specific functional form is assumed for the PDFs or whether a non-parametric approach is adopted. Although the former approach has been widely used in the literature, non-parametric approaches based on Bayesian inference have been successfully applied to the problem of PDF determination, albeit in a limited scenario [DelDebbio:2021whr, Candido:2024hjt, Medrano:2025cmg]. Bayesian-based approaches are promising tools that ensure a rigorous framework where prior information and assumptions are spelled out explicitly. Yet, a global PDF determination based on these methods has not yet been attempted, and the impact of the prior needs to be carefully studied in these frameworks.
@@
On the other hand, state-of-the-art PDF determinations rely on parametric approaches, where a specific functional form is assumed for the PDFs at a given initial scale $Q_{0}$. These functions are typically chosen to be flexible enough to capture the main features of the PDFs, while their internal parameters are optimised to reproduce the experimental data. Several groups [NNPDF:2021njg, Ablat:2024hbm, Bailey:2020ooq, Alekhin:2017kpj] have set the standard for PDF determinations through continuous refinement of their global fits as new data and theoretical advances become available, with an increasing emphasis on uncertainty quantification. Although these determinations have been shown to perform well on a wide range of new experimental data [Chiefa:2025loi], the different methodological frameworks adopted by the various groups lead to PDF sets whose differences are yet to be fully understood [Harland-Lang:2024kvt, PDF4LHCWorkingGroup:2022cjn]. These differences become significantly visible when considering parameter determinations that are particularly sensitive to the choice of the PDF set, both on the central values and, most importantly, on the associated uncertainties (see Refs. [ATLAS:2023lhg, CMS:2024ony, ATLAS:2023fsi] for some recent examples).
@@
In this work, we build upon the work of Refs. [DelDebbio:2021whr, Candido:2024hjt], which aims at providing a sound statistical framework for PDF determination, with all underlying assumptions clearly stated. We focus on the NNPDF methodology [NNPDF:2021njg], which pioneered the use of ML tools in the context of PDF determinations and has been validated through extensive studies over the years [DelDebbio:2021whr, Barontini:2025lnl, Cruz-Martinez:2021rgy]. It combines a Monte Carlo sampling of the experimental data and a feed-forward neural network parametrization of the PDFs. We adopt a simplified framework to analyse the training process, aiming at providing a quantitative description of its key aspects, and making transparent the assumptions that are often implicitly embedded in the fitting procedure.
@@
We demonstrate that the training dynamics of a neural network can be fully reformulated in functional space, leading to an interpretable description of the learning process. We show that the training dynamics is dictated by the Neural Tangent Kernel (NTK) [jacot2018neural], which encodes and factorises the dependence on the architecture and the parameters of the neural network. Similar approaches leveraging NTK properties have been explored in other contexts [tovey2025collective]. In fact, the spectral properties of the NTK provide a powerful lens through which we can understand the learning process: only the directions that are orthogonal to the kernel of the NTK are actually learned in the training process. At initialisation, the NTK is characterised by a wide spectrum of eigenvalues, with only a few large eigenvalues being significantly different from zero. Even though the NNs span a very broad functional space, the training explores a much smaller subspace. During the training process, the hierarchy in the NTK spectrum is preserved, but eigenvalues that were initially subleading, or zero, grow in magnitude. Since the only directions that contribute to the learning process are those associated to the non-zero eigenvalues, with the actual value of the eigenvalue setting the learning speed along the corresponding eigenvector direction, the growth of some eigenvalues indicates that new features in the functional space emerge during training and that the network thus becomes capable of representing more complex functions. The space of functions explored in the training process is therefore dynamically determined during the training itself, exploiting the flexibility of the parametrization in order to explore multiple functional forms and select the preferred one based on inference from data rather than a priori decisions.
@@
Another key result of this work is that, after an initial transient phase where the NTK evolves significantly, the training process enters a second regime where the NTK becomes approximately constant. This regime is often referred to as lazy training in the Machine Learning literature [jacot2018neural], and it has important implications for the training dynamics. In this regime, we show that the training process can be described analytically, allowing us to obtain a single and clean closed-form expression for the output of the network at any training time $t$. The main result of this analysis is summarised in Eq. (56) which we report here,
@@ 2 Neural Networks and PDFs
In the following, we prepare the ground for the study of the training dynamics of neural networks used in the NNPDF framework. We start by briefly presenting the inverse problem of PDF determination using data depending linearly on the PDFs, setting the notation and introducing the statistical vocabulary used in the rest of this study. We then discuss some statistical aspects of the neural networks at initialisation, which will help us understand the implications in the training process. These properties, derived in the large-width limit [lee2019wide, jacot2018neural], are analysed for the specific architecture used in the NNPDF methodology. An exhaustive and detailed review of wide-network properties is beyond the scope of this work, and the reader is encouraged to refer to Ref. [Roberts:2021fes] for a comprehensive review.
@@ 2.1 The 1-dimensional regression problem of PDFs
The extraction of PDFs from experimental data is a classic example of an inverse problem, namely the reconstruction of a function $f(x)$ from a finite set of data points $Y_{I}$, where the index $I=1,\ldots,N_{\mathrm{dat}}$ [^1] When omitting the data index $I$, we will always assume $Y\in\mathbb{R}^{N_{\mathrm{dat}}}$.. In particular, for this study, we will focus on DIS data, which depend linearly on the function $f(x)$. The theoretical prediction for the data point $Y_{I}$ is given by
@@
Attempting to determine a function $f$ in an infinite dimensional space of solutions using a finite set of data is inherently ill-posed. The solution inevitably depends on assumptions and prior knowledge – conscious or not – introduced to regularise the problem. Different methodologies, based either on non-parametric methods or parametric regression, have been proposed to address these challenges, yielding increasingly precise PDFs. Yet, despite the longstanding effort to provide robust uncertainty quantification and establish the relationships between different methodologies and their solutions, some discrepancies remain unresolved, see, e.g., [PDF4LHCWorkingGroup:2022cjn]. Understanding such differences between the various approaches is thus crucial for precision physics.
Following the ideas highlighted in Refs. [DelDebbio:2021whr, Candido:2024hjt], the solution of the inverse problem is conveniently phrased in a Bayesian framework. The functions $f_{i}$ are promoted to stochastic processes; for any grid of points $x_{\alpha}$, $\alpha=1,\ldots,N_{\mathrm{grid}}$, the vector $f_{i\alpha}=f_{i}(x_{\alpha})$ is a vector of $N_{f}\times N_{\mathrm{grid}}$ stochastic variables, for which we introduce a prior distribution $p(f)$  [^2] Following the same convention used for the data, when omitting the grid index $\alpha$, and/or the flavor index $i$, we will always refer to a vector $f\in\mathbb{R}^{N_{f}\times N_{\mathrm{grid}}}$.. In this perspective, any fitting procedure is interpreted as a recipe that yields the posterior distribution $\tilde{p}(f)=p(f|Y)$. In this study, following the NNPDF methodology, probability distributions are represented by ensembles of i.i.d. neural network replicas. So, for instance, the prior distribution $p(f)$ is described by an ensemble
@@
The prior distribution $p(f)$ is defined by initializing a set of neural networks (NNs) replicas using a Glorot normal initializer [glorot2010understanding]. The result of this initialisation is discussed below in Sec. 2.2.
In order to account for the experimental uncertainties and propagate them to the fitted PDFs, the NNPDF collaboration uses Monte Carlo replicas. For each replica, labelled by the index $k$, a new set of data $Y^{(k)}$ is generated from an $N_{\mathrm{dat}}$ dimensional Gaussian distribution centred at the experimental central value $Y$, with the covariance given by the experimental covariance matrix $C_{Y}$,
@@ 2.2 Neural Networks at Initialisation
When initializing a neural network, the weights and biases – which we denote collectively as the parameters of the network – are drawn from some probability distribution. In the NNPDF formalism, the set of network parameters at initialisation for each replica is an instance of i.i.d. stochastic variables. More importantly, the probability distribution of the network parameters induces a probability distribution for the output of the neural networks at initialisation. It is well known that the probability distribution of these outputs becomes approximately Gaussian when the size of the hidden layers is increased  [Roberts:2021fes]. We call this limit the large-network limit.
As detailed in Ref. [NNPDF:2021njg], the NNs used for the NNPDF fit have a 2-25-20-8 architecture, a $\tanh$ activation function, and are initialized using a Glorot normal distribution [glorot2010understanding]. The preactivation function of a neuron is denoted as $\phi^{(\ell)}_{i,\alpha}=\phi^{(\ell)}_{i}(x_{\alpha})$, where $\ell=0,\ldots,L$, denotes the layer of the neuron, and, for each $\ell$, $i=1,\ldots,n_{\ell}$ identifies the neuron within the layer [^3] We refer to $i$ as the neuron index.. Furthermore, $x_{\alpha}$ is the input to the NN, i.e., a point in the interval $[0,1]$. A grid of $N_{\mathrm{grid}}=50$ points in $x$ is used to compute observables in the NNPDF formalism and in this work we focus on the values of $f$ at those values of $x_{\alpha}$, where the index $\alpha=1,\ldots,N_{\mathrm{grid}}$ labels the points on the grid. For completeness, we list the values of $x_{\alpha}$ in Tab. 1.
@@
The PDFs in the so-called evolution basis are parametrized by the preactivation functions of the output layer $L$, $x_{\alpha}f_{i}(x_{\alpha})=\phi^{(L)}_{i,\alpha}$, where the neuron index on the last layer, $i=1,\ldots,8$, labels the flavors [^4] For simplicity, we ignore the preprocessing function $x^{-\alpha_{i}}(1-x)^{\beta_{i}}$ that is currently used in the NNPDF fits. While the preprocessing may be useful in speeding the training it does not affect the current discussion.. The input layer is identified by $\ell=0$ and the activation function for that specific layer is the identity, so that
@@
This duality between parameter-space and function-space provides a powerful framework to study the behaviour of an ensemble of NNs, and in particular the symmetry properties of the distribution $p(\phi^{(\ell)})$ (see, e.g., Ref. [Maiti:2021fpy]). Working in parameter space, i.e., computing the expectation values of correlators of fields as integrals over the NN parameters, one can readily show that
@@
so that the probability distribution is fully determined by the couplings $\gamma^{(\ell)}$. [^5] We have denoted all couplings by $\gamma^{{(\ell)}}$. Different couplings are identified by the number of indices, so that $\gamma^{(\ell)}_{\alpha_{1}\alpha_{2}}$ is a two-point coupling, $\gamma^{(\ell)}_{\alpha_{1}\alpha_{2},\alpha_{3}\alpha_{4}}$ is a four-point coupling, etc. In Eq. (15), we have factored out inverse powers of $n_{\ell-1}$ for each coupling. With this convention, and with the scaling of the parameters variances in Eq. (9), the couplings in the action are all $O(1)$ in the limit where $n_{\ell}\to\infty$. As a consequence, the probability distribution at initialisation is a multidimensional Gaussian at leading order – i.e., $\mathcal{O}(1)$ – in $1/n_{\ell}$, with quartic corrections that are $O(1/n_{\ell})$, while higher powers of the invariant bilinear are suppressed by higher powers of the width of the layer. This power counting defines an effective field theory, where deviations from Gaussianity can be computed in perturbation theory to any given order in $1/n_{\ell}$, see, e.g.  Ref. [Roberts:2021fes, Chiefa:2026TBA] for a detailed presentation of these ideas. While the actual calculations become rapidly cumbersome, the conceptual framework is straightforward.
@@ 3 Training Dynamics and the Neural Tangent Kernel
In practice, minimization is performed using gradient methods where the direction towards the minimum is defined by the gradient of the loss function. These methods are usually improved by including, for instance, stochasticity and information on previous iterations [amary1993sdg, Nadam2016, kingma2014adam]. A detailed overview of these extended gradient methods is beyond the scope of this work. In the context of PDF determinations, the NNPDF collaboration makes intensive use of these tools and the reader is encouraged to refer to Ref. [NNPDF:2021njg] for an extensive discussion.
@@
Finally, we emphasise that the results in this section, while obtained having in mind neural networks, apply to any fixed parametrization, including fixed functional forms [Ablat:2024hbm, Bailey:2020ooq, Alekhin:2017kpj] or kernels [Costantini:2025wxp].
@@ 3.1 Training in Functional Space
For analytical tractability, GD is described as a continuous flow of the parameters $\theta$ in training time $t$ along the negative gradient of the loss function $\mathcal{L}$. For sufficiently small learning rates $\eta$, this continuous flow approximates the discrete GD trajectory in parameter space, as extensively discussed in Ref. [barrett2022igr]. The continuous Gradient Flow (GF) is then given by
@@
where $\theta_{t,\mu}$ and $\mathcal{L}_{t}$ identify respectively the parameters and the loss function at training time $t$. We distinguish between the continuous training time $t$ and the discrete epochs of GD, the latter denoted using the capital letter $T$. The two are related through the learning rate, $t=\eta T$, and in the following we will use them interchangeably.
@@
where we have defined the Neural Tangent Kernel [jacot2018neural]
@@
As it will become clearer later, the NTK provides a powerful framework for understanding neural network dynamics during training. Originally developed by Jacot et al. [jacot2018neural] to analyse infinite-width feed-forward networks, the NTK theory has since been extended to diverse architectures including convolutional networks [arora2019exact] and recurrent networks [alemohammad2021recurrent]. This theoretical framework has proven invaluable for characterizing learning dynamics and generalization properties across various network designs. We will see in the following and in Sec. 4 how the NTK can also provide useful insights in the context of PDF fitting.
@@
Before moving to the next subsection, a few comments are in order. First, although derived in the context of neural networks, these equations do not refer to a specific parametrization. Indeed, these remain valid even when an explicit functional form is chosen to parametrize the PDFs, as in Refs. [Ablat:2024hbm, Bailey:2020ooq, Costantini:2025wxp]. Second, it is interesting to observe that the flow equation, Eq. (34), depends on two matrices, $\Theta$ and $M$. The former encodes the model dependence, while the latter contains the physical information. The interplay between these two matrices is crucial for understanding the training dynamics, as discussed in Sec. 3.2.3. Finally, the NTK derived in Eq. (32) is inherently time-dependent in a complex way, which precludes any attempt at integrating Eq. (34) analytically. We come back to this point in Sec. 4.1.1, after discussing the properties of the NTK during training.

### v2

@@ 1 Introduction
The complex nature of inverse problems has prompted the development of sophisticated statistical methods and tools to tackle them. In general, PDF determinations can be broadly classified into two main categories, depending on whether a specific functional form is assumed for the PDFs or whether a non-parametric approach is adopted. Although the former approach has been widely used in the literature, non-parametric approaches based on Bayesian inference have been successfully applied to the problem of PDF determination, albeit in a limited scenario [21, 14, 32]. Bayesian-based approaches are promising tools that ensure a rigorous framework where prior information and assumptions are spelled out explicitly. Yet, a global PDF determination based on these methods has not yet been attempted, and the impact of the prior needs to be carefully studied in these frameworks.
@@
On the other hand, state-of-the-art PDF determinations rely on parametric approaches, where a specific functional form is assumed for the PDFs at a given initial scale $Q_{0}$. These functions are typically chosen to be flexible enough to capture the main features of the PDFs, while their internal parameters are optimised to reproduce the experimental data. Several groups [9, 2, 7, 3] have set the standard for PDF determinations through continuous refinement of their global fits as new data and theoretical advances become available, with an increasing emphasis on uncertainty quantification. Although these determinations have been shown to perform well on a wide range of new experimental data [15], the different methodological frameworks adopted by the various groups lead to PDF sets whose differences are yet to be fully understood [25, 10]. These differences become significantly visible when considering parameter determinations that are particularly sensitive to the choice of the PDF set, both on the central values and, most importantly, on the associated uncertainties (see Refs. [1, 26, 27] for some recent examples).
@@
In this work, we build upon the work of Refs. [21, 14], which aims at providing a sound statistical framework for PDF determination, with all underlying assumptions clearly stated. We focus on the NNPDF methodology [9], which pioneered the use of ML tools in the context of PDF determinations and has been validated through extensive studies over the years [21, 11, 20]. It combines a Monte Carlo sampling of the experimental data and a feed-forward neural network parametrization of the PDFs. We adopt a simplified framework to analyse the training process, aiming at providing a quantitative description of its key aspects, and making transparent the assumptions that are often implicitly embedded in the fitting procedure.
@@
We demonstrate that the training dynamics of a neural network can be fully reformulated in functional space, leading to an interpretable description of the learning process. We show that the training dynamics is dictated by the Neural Tangent Kernel (NTK) [28], which encodes and factorises the dependence on the architecture and the parameters of the neural network. Similar approaches leveraging NTK properties have been explored in other contexts [34]. In fact, the spectral properties of the NTK provide a powerful lens through which we can understand the learning process: only the directions that are orthogonal to the kernel of the NTK are actually learned in the training process. At initialisation, the NTK is characterised by a wide spectrum of eigenvalues, with only a few large eigenvalues being significantly different from zero. Even though the NNs span a very broad functional space, the training explores a much smaller subspace. During the training process, the hierarchy in the NTK spectrum is preserved, but eigenvalues that were initially subleading, or zero, grow in magnitude. Since the only directions that contribute to the learning process are those associated to the non-zero eigenvalues, with the actual value of the eigenvalue setting the learning speed along the corresponding eigenvector direction, the growth of some eigenvalues indicates that new features in the functional space emerge during training and that the network thus becomes capable of representing more complex functions. The space of functions explored in the training process is therefore dynamically determined during the training itself, exploiting the flexibility of the parametrization in order to explore multiple functional forms and select the preferred one based on inference from data rather than a priori decisions.
@@
Another key result of this work is that, after an initial transient phase where the NTK evolves significantly, the training process enters a second regime where the NTK becomes approximately constant. This regime is often referred to as lazy training in the Machine Learning literature [28], and it has important implications for the training dynamics. In this regime, we show that the training process can be described analytically, allowing us to obtain a single and clean closed-form expression for the output of the network at any training time $t$. The main result of this analysis is summarised in Eq. (57) which we report here,
@@ 2 Neural Networks and PDFs
In the following, we prepare the ground for the study of the training dynamics of neural networks used in the NNPDF framework. We start by briefly presenting the inverse problem of PDF determination using data depending linearly on the PDFs, setting the notation and introducing the statistical vocabulary used in the rest of this study. We then discuss some statistical aspects of the neural networks at initialisation, which will help us understand the implications in the training process. These properties, derived in the large-width limit [30, 28], are analysed for the specific architecture used in the NNPDF methodology. An exhaustive and detailed review of wide-network properties is beyond the scope of this work, and the reader is encouraged to refer to Ref. [33] for a comprehensive review.
@@ 2.1 The 1-dimensional regression problem of PDFs
The extraction of PDFs from experimental data is a classic example of an inverse problem, namely the reconstruction of a function $f(x)$ from a finite set of data points $Y_{I}$, where the index $I=1,\ldots,N_{\mathrm{dat}}$. [^1] When omitting the data index $I$, we will always assume $Y\in\mathbb{R}^{N_{\mathrm{dat}}}$. In particular, for this study, we will focus on DIS data, which depend linearly on the function $f(x)$. The theoretical prediction for the data point $Y_{I}$ is given by
@@
Attempting to determine a function $f$ in an infinite dimensional space of solutions using a finite set of data is inherently ill-posed. The solution inevitably depends on assumptions and prior knowledge – conscious or not – introduced to regularise the problem. Different methodologies, based either on non-parametric methods or parametric regression, have been proposed to address these challenges, yielding increasingly precise PDFs. Yet, despite the longstanding effort to provide robust uncertainty quantification and establish the relationships between different methodologies and their solutions, some discrepancies remain unresolved, see, e.g., [10]. Understanding such differences between the various approaches is thus crucial for precision physics.
Following the ideas highlighted in Refs. [21, 14], the solution of the inverse problem is conveniently phrased in a Bayesian framework. The functions $f_{i}$ are promoted to stochastic processes; for any grid of points $x_{\alpha}$, $\alpha=1,\ldots,N_{\mathrm{grid}}$, the vector $f_{i\alpha}=f_{i}(x_{\alpha})$ is a vector of $N_{f}\times N_{\mathrm{grid}}$ stochastic variables, for which we introduce a prior distribution p(f)p(f). [^2] Following the same convention used for the data, when omitting the grid index $\alpha$, and/or the flavor index $i$, we will always refer to a vector $f\in\mathbb{R}^{N_{f}\times N_{\mathrm{grid}}}$. In this perspective, any fitting procedure is interpreted as a recipe that yields the posterior distribution $\tilde{p}(f)=p(f|Y)$. In this study, following the NNPDF methodology, probability distributions are represented by ensembles of i.i.d. neural network replicas. So, for instance, the prior distribution $p(f)$ is described by an ensemble
@@
The prior distribution $p(f)$ is defined by initializing a set of neural networks (NNs) replicas using a Glorot normal initializer [24]. The result of this initialisation is discussed below in Sec. 2.2.
In order to account for the experimental uncertainties and propagate them to the fitted PDFs, the NNPDF collaboration uses Monte Carlo replicas [17]. For each replica, labelled by the index $k$, a new set of data $Y^{(k)}$ is generated from an $N_{\mathrm{dat}}$ dimensional Gaussian distribution centred at the experimental central value $Y$, with the covariance given by the experimental covariance matrix $C_{Y}$,
@@ 2.2 Neural Networks at Initialisation
When initializing a neural network, the weights and biases – which we denote collectively as the parameters of the network – are drawn from some probability distribution. In the NNPDF formalism, the set of network parameters at initialisation for each replica is an instance of i.i.d. stochastic variables. More importantly, the probability distribution of the network parameters induces a probability distribution for the output of the neural networks at initialisation. It is well known that the probability distribution of these outputs becomes approximately Gaussian when the size of the hidden layers is increased  [33]. We call this limit the large-network limit.
As detailed in Ref. [9], the NNs used for the NNPDF fit have a 2-25-20-8 architecture, a $\tanh$ activation function, and are initialized using a Glorot normal distribution [24]. The preactivation function of a neuron is denoted as $\phi^{(\ell)}_{i,\alpha}=\phi^{(\ell)}_{i}(x_{\alpha})$, where $\ell=0,\ldots,L$, denotes the layer of the neuron, and, for each $\ell$, $i=1,\ldots,n_{\ell}$ identifies the neuron within the layer. [^3] We refer to $i$ as the neuron index. Furthermore, $x_{\alpha}$ is the input to the NN, i.e., a point in the interval $[0,1]$. A grid of $N_{\mathrm{grid}}=50$ points in $x$ is used to compute observables in the NNPDF formalism and in this work we focus on the values of $f$ at those values of $x_{\alpha}$, where the index $\alpha=1,\ldots,N_{\mathrm{grid}}$ labels the points on the grid. For completeness, we list the values of $x_{\alpha}$ in Tab. 1.
@@
The PDFs in the so-called evolution basis are parametrized by the preactivation functions of the output layer $L$, $x_{\alpha}f_{i}(x_{\alpha})=\phi^{(L)}_{i,\alpha}$, where the neuron index on the last layer, $i=1,\ldots,8$, labels the flavors. [^4] For simplicity, we ignore the preprocessing function $x^{-\alpha_{i}}(1-x)^{\beta_{i}}$ that is currently used in the NNPDF fits. While the preprocessing may be useful in speeding the training it does not affect the current discussion. The input layer is identified by $\ell=0$ and the activation function for that specific layer is the identity, so that
@@
This duality between parameter-space and function-space provides a powerful framework to study the behaviour of an ensemble of NNs, and in particular the symmetry properties of the distribution $p(\phi^{(\ell)})$ (see, e.g., Ref. [31]). Working in parameter space, i.e., computing the expectation values of correlators of fields as integrals over the NN parameters, one can readily show that
@@
so that the probability distribution is fully determined by the couplings $\gamma^{(\ell)}$. [^5] We have denoted all couplings by $\gamma^{{(\ell)}}$. Different couplings are identified by the number of indices, so that $\gamma^{(\ell)}_{\alpha_{1}\alpha_{2}}$ is a two-point coupling, $\gamma^{(\ell)}_{\alpha_{1}\alpha_{2},\alpha_{3}\alpha_{4}}$ is a four-point coupling, etc. In Eq. (15), we have factored out inverse powers of $n_{\ell-1}$ for each coupling. With this convention, and with the scaling of the parameters variances in Eq. (9), the couplings in the action are all $O(1)$ in the limit where $n_{\ell}\to\infty$. As a consequence, the probability distribution at initialisation is a multidimensional Gaussian at leading order – i.e., $\mathcal{O}(1)$ – in $1/n_{\ell}$, with quartic corrections that are $O(1/n_{\ell})$, while higher powers of the invariant bilinear are suppressed by higher powers of the width of the layer. This power counting defines an effective field theory, where deviations from Gaussianity can be computed in perturbation theory to any given order in $1/n_{\ell}$, see, e.g.  Ref. [33, 16] for a detailed presentation of these ideas. While the actual calculations become rapidly cumbersome, the conceptual framework is straightforward.
@@ 3 Training Dynamics and the Neural Tangent Kernel
In practice, minimization is performed using gradient methods where the direction towards the minimum is defined by the gradient of the loss function. These methods are usually improved by including, for instance, stochasticity and information on previous iterations [5, 22, 29]. A detailed overview of these extended gradient methods is beyond the scope of this work. In the context of PDF determinations, the NNPDF collaboration makes intensive use of these tools and the reader is encouraged to refer to Ref. [9] for an extensive discussion.
@@
Finally, we emphasise that the results in this section, while obtained having in mind neural networks, apply to any fixed parametrization, including fixed functional forms [2, 7, 3] or kernels [19].
@@ 3.1 Training in Functional Space
For analytical tractability, GD is described as a continuous flow of the parameters $\theta$ in training time $t$ along the negative gradient of the loss function $\mathcal{L}$. For sufficiently small learning rates $\eta$, this continuous flow approximates the discrete GD trajectory in parameter space, as extensively discussed in Ref. [12]. The continuous Gradient Flow (GF) is then given by
@@
where $\theta_{t,\mu}$ and $\mathcal{L}_{t}$ identify respectively the parameters and the loss function at training time $t$. We distinguish between the continuous training time $t$ and the discrete epochs of GD, the latter denoted using the capital letter $T$. The two are related through the learning rate, $t=\eta T$ with $\eta=10^{-5}$, and in the following we will use them interchangeably.
@@
where we have defined the Neural Tangent Kernel [28]
@@
As it will become clearer later, the NTK provides a powerful framework for understanding neural network dynamics during training. Originally developed by Jacot et al. [28] to analyse infinite-width feed-forward networks, the NTK theory has since been extended to diverse architectures including convolutional networks [6] and recurrent networks [4]. This theoretical framework has proven invaluable for characterizing learning dynamics and generalization properties across various network designs. We will see in the following and in Sec. 4 how the NTK can also provide useful insights in the context of PDF fitting.
@@
Before moving to the next subsection, a few comments are in order. First, although derived in the context of neural networks, these equations do not refer to a specific parametrization. Indeed, these remain valid even when an explicit functional form is chosen to parametrize the PDFs, as in Refs. [2, 7, 19]. Second, it is interesting to observe that the flow equation, Eq. (34), depends on two matrices, $\Theta$ and $M$. The former encodes the model dependence, while the latter contains the physical information. The interplay between these two matrices is crucial for understanding the training dynamics, as discussed in Sec. 3.2.3. Finally, the NTK derived in Eq. (32) is inherently time-dependent in a complex way, which precludes any attempt at integrating Eq. (34) analytically. We come back to this point in Sec. 4.1.1, after discussing the properties of the NTK during training.